Given that P(A)=0.7,P(B)= and P (A and B)=0.35


1
Expert's answer
2021-06-16T05:35:18-0400

Given "P(A)=0.7, P(B)=0.6," and "P(A\cap B)=0.35."

A.


"P(B\cap A')=P(B)-P(A\cap B)"

"=0.6-0.35=0.25"

B.


"P(B')=1-P(B)"

"=1-0.6=0.4"

C.


"P(A)P(B)=0.7(0.6)=0.42\not=0.35=P(A\cap B)"

Therefore "A" and "B" are dependent events.


D.


"P(A\cup B)=P(A)+P(B)-P(A\cap B)"

"=0.7+0.6-0.35=0.95"

E.


"P(B|A)=\dfrac{P(A\cap B)}{P(A)}"

"=\dfrac{0.35}{0.7}=0.5"

F.


"P(A\cap B)=0.35\not=0"

Therefore "A" and "B" are not mutually exclusive dependent events. e